# No Images, No Problem: Retaining Knowledge in Continual VQA with Questions-Only Memory

### v1

@@ Abstract
Continual Learning in Visual Question Answering (VQACL) requires models to learn new visual-linguistic tasks (plasticity) while retaining knowledge from previous tasks (stability). The multimodal nature of VQACL presents unique challenges, requiring models to balance stability across visual and textual domains while maintaining plasticity to adapt to novel objects and reasoning tasks. Existing methods, predominantly designed for unimodal tasks, often struggle to balance these demands effectively. In this work, we introduce QUestion-only replay with Attention Distillation (QUAD), a novel approach for VQACL that leverages only past task questions for regularization, eliminating the need to store visual data and addressing both memory and privacy concerns. QUAD achieves stability by introducing a Question-only Replay mechanism that selectively uses questions from previous tasks to prevent overfitting to the current task’s answer space, thereby mitigating the out-of-answer-set problem. Complementing this, we propose Attention Consistency Distillation, which uniquely enforces both intra-modal and inter-modal attention consistency across tasks, preserving essential visual-linguistic associations. Extensive experiments on VQAv2 and NExT-QA demonstrate that QUAD significantly outperforms state-of-the-art methods, achieving robust performance in continual VQA [^1] Work done during an internship at Computer Vision Center (CVC), Universitat Autonoma de Barcelona.. Code is available at: https://github.com/IemProg/QUAD
@@ 1 Introduction
Continual learning (CL) seeks to enable models to incorporate new information while retaining previously learnt knowledge, thereby addressing the problem of catastrophic forgetting (CF) [48, 49]. This capability is essential in dynamic real-world settings, where models must continuously adapt to evolving data while preserving previously acquired knowledge [85]. While significant advances have been made in CL, research has predominantly focused on unimodal tasks, particularly image classification [70, 69, 33, 22, 47, 87]. However, real-world applications often demand more sophisticated multimodal learning approaches, especially for tasks that integrate visual and textual information. Visual Question Answering (VQA) stands as a representative multimodal task, requiring models to simultaneously process and reason over both images and natural language. For instance, answering questions like “What color is the car on the left?” or “How many trucks are visible?” demands both object recognition from the image and an understanding of the linguistic query structure [85, 60, 78, 8, 72].
@@
The emerging field of Visual Question Answering Continual Learning (VQACL) [85] presents unique challenges at the intersection of continual learning and multimodal reasoning. VQACL requires maintaining a balance between stability—the preservation of past knowledge—and plasticity—the ability to learn new information—across both visual and linguistic modalities [85, 3]. This dual-modality requirement, combined with the need for generalisation ability, introduces additional complexity in learning dynamics. Models must not only retain their understanding of previously encountered visual elements and language patterns but also generalise to unseen objects and novel question formats. For example, a model that masters counting vehicles should be able to transfer this counting capability to novel object categories, such as bicycles [23].
@@
Memory-based methods in CL [7, 6, 67, 85] effectively mitigate catastrophic forgetting through storing and replaying exemplar data. In the VQACL setting, these approaches store complete image-question pairs, necessitating significant memory resources—often thousands per task (e.g., 5000 samples in VQACL [85]). This extensive storage of visual data raises two critical challenges. First, it imposes substantial computational and storage burdens, which can be infeasible in resource-constrained real-world applications. Second, and more importantly, it raises profound privacy concerns. Visual data often contains sensitive and personally identifiable information, especially in domains like healthcare, finance, and surveillance [64, 42], where stringent regulations such as GDPR [19, 81] govern data usage and storage. In contrast, textual data (the questions) is typically non-identifiable and generic, inherently posing less privacy risk. Memory-free methods [35, 1, 83, 33], though eliminating storage and privacy concerns by avoiding data retention entirely, exhibit suboptimal performance in multimodal settings like VQACL. This dichotomy between memory-rehearsal and memory-free approaches raises an intriguing questions: Is it truly necessary to store visual data? Could storing only past questions suffice to mitigate forgetting? To explore these questions, we propose a novel intermediate setting: VQACL with Question-only Rehearsal (VQACL-QR) (Fig. 1). In this framework, only questions from past tasks are stored, eliminating the need for visual data storage. It provides a practical, privacy-preserving solution for continual VQA setting.
@@
To address VQACL-QR , we propose QUestion-only replay with Attention Distillation (QUAD), a novel distillation method to balance stability and plasticity using only past questions. Our approach introduces two key contributions, each targeting specific challenges in continual VQA. First, we introduce a Question-only Replay mechanism, which utilises only questions from past tasks to regularise the current model with a strategic question selection process that directly tackles the out-of-answer-set problem – model tend to overfit to the current task’s answer space, leading to incorrect responses for previous tasks’ questions (see Fig. 1). Second, we propose attention consistency distillation, a novel attention distillation strategy that preserves the model’s attention patterns across tasks. Our distillation strategy enforces both intra-modal consistency (language-to-language, visual-to-visual) and inter-modal consistency (language-to-visual), ensuring that the model consistently focuses on relevant regions while adapting to new information.
Through extensive experiments on standard VQACL benchmarks (VQAv2, and NExT-QA datasets), we demonstrate that QUAD achieves state-of-the-art performance, surpassing both memory-free approaches and traditional rehearsal methods that store complete image-question pairs. Our findings suggest that storing only questions is sufficient to mitigate forgetting, validating the practicality of the Question-only Rehearsal (VQACL-QR) setting.
@@ 2 Related work
Visual Question Answering (VQA) involves answering natural language questions by interpreting visual content [3, 51, 38]. Recent approaches leverage vision-language models (VLMs) built on transformer architectures [16, 61, 76, 37, 8] alongside pre-trained language models [78, 80]. For instance, Cho et al. [9] proposed a generative transformer for VQA that analyses visual elements in conjunction with textual questions. Many methods aim to improve generalisation capabilities by enhancing compositionality, which is essential for cognitive reasoning [30, 36]. For instance, Johnson et al. [28] explored the composition of visual attributes by creating a dataset designed for compositional reasoning, while Whitehead et al. [73] used contrastive learning to enhance compositionality, disentangling reasoning skills from visual concepts. Although these approaches make strides toward compositionality, the implicit decomposition may impact generalisation, and constructing effective contrastive samples remains complex.
Continual Learning (CL) aims to develop frameworks capable of incrementally learning from sequentially arriving datasets. This is a fundamental challenge for many deep learning methods due to catastrophic forgetting (CF) [48]. CL methods can be divided into different categories. While knowledge distillation-based methods constrain mappings between incremental models to prevent forgetting [39, 57, 27, 46, 13, 29], optimisation-based techniques modify the gradient updates during training to minimise interference with previously learnt tasks to mitigate forgetting [44, 5, 59, 79], and representation-based methods focus on learning robust and adaptable feature representations [18, 17, 15, 14]. Recently we have also witnessed the uprisal of prompt-based methods like [71, 70, 68, 62], which utilise visual prompts [43] with pre-trained transformers for CL scenarios. Most of these methods target Class-Incremental Learning (CIL) in unimodal settings; as such, they cannot simultaneously tackle multimodal continuous data and ignore the essential composition generalisation issue for VQA [85, 52]. Our work falls within the distillation-based category, as QUAD enables distillation with only past questions to reduce CF for continual VQA.
Continual VQA. Recent research has increasingly focused on multimodal CL in the context of VQA [10, 24, 52, 63]. While studies like Greco et al. [24] have examined forgetting dynamics in VQA, they often focus on a limited set of tasks and overlook the potential of pre-trained models to mitigate forgetting [50]. The VQACL benchmark [85] evaluated conventional continual learning approaches such as  [7, 6, 67, 85]. Some works have explored specific aspects of continual VQA, such as question types [24], domain adaptation [86], and compositionality [85]. Our work introduces a distillation method for continual VQA while remaining memory-efficient, privacy-conscious, and without any stored prototypes as in the VQACL method [^2] VQACL represents both the setting and the approach for continual learning in Visual Question Answering, as described in  [85]. [85].
@@ 3.1 Setting Overview
In VQACL, the sequence of tasks is organized as a series of $L$ macro-tasks, each comprising $K$ sub-tasks, resulting in a total of $T=L\times K$ tasks. Each macro-task is designed to develop specific reasoning skills such as counting, color identification, or object recognition (i.e linguistic task). For example, in a counting task, the model primarily engages with questions like “How many objects are there?” or “What number is shown?”.
@@
Figure 2: Overview of QUAD. Our method consists of three key components designed to maintain stability and plasticity in the VQACL setting. (1) Questions-Only Memory ($\mathcal{M}$) stores questions from past tasks, without visual data. (2) Pseudo-label Distillation ($\mathcal{L}_{\text{PL}}$) leverages answers generated by the previous model $\theta^{t-1}$ for new image-question pairs, guiding the current model $\theta^{t}$ to retain past knowledge, and mitigate the out-of-answer-set problem. (3) Attention Consistency Distillation ($\mathcal{L}_{\text{Att}}$) aligns the self-attention maps between $\theta^{t-1}$ and $\theta^{t}$ to maintain focus on relevant visual-linguistic relationships. The task-specific loss ($\mathcal{L}_{\text{CE}}$) is applied solely to current task samples, promoting adaptation to new data.
@@ 3.2 Overview
We propose QUAD, a novel approach for VQACL-QR that eliminates the need to store images from previous tasks by using only past task questions (see Fig. 2). Following several works in continual learning [40, 11], we adopt a regularisation framework where the total learning objective $\mathcal{L}_{\text{VQACL}}$ consists of two main components:
$\mathcal{L}_{\text{VQACL}}=\mathcal{L}_{\text{Plasticity}}+\lambda\mathcal{L}_{\text{Stability}}.$ (1)
The first, the plasticity term $\mathcal{L}_{\text{Plasticity}}$, guides the model’s adaptation to the current task $\mathcal{T}^{t}$, and $\lambda>0$ is a weighting coefficient adjusting the balance between plasticity and stability. Following common practice in VQA [85, 56, 2], we implement this loss using cross-entropy to compare the network’s prediction for an input image-question pair $(x^{t},q^{t})$ with the corresponding annotated answer $y^{t}$:
@@
The second loss component, the stability term $\mathcal{L}_{\text{Stability}}$, helps prevent forgetting. In the standard VQACL setting, where images from past tasks can be stored in the memory $\mathcal{M}$, this term can be implemented similarly to the plasticity term, using cross-entropy and averaging the loss over triplets ${}{(x^{m},q^{m},y^{m})\in\mathcal{M}}$. In the VQACL-QR setting, storage of images $x^{m}$ from past tasks in the memory $\mathcal{M}$ is prohibited. To address this limitation, we propose a novel $\mathcal{L}_{\text{Stability}}$ loss tailored for VQACL-QR. This loss combines two complementary strategies—pseudo-labeling $\mathcal{L}_{\text{PL}}$ and attention distillation $\mathcal{L}_{\text{Att}}$— to effectively compensate for the absence of past task images, ensuring robust knowledge retention while adhering to the constraints of the VQACL-QR framework. In our case, the stability term can be written as:
$\mathcal{L}_{\text{Stability}}=\mathcal{L}_{\text{PL}}+\mathcal{L}_{\text{Att}},$ (3)
In the following, we explain how we implement these two losses without using any images from past tasks.
3.3 Pseudo-Labeling for Question-only Replay
@@
To address knowledge retention in continual VQA, we propose a pseudo-labeling approach that leverages image-question pairs formed by combining current task images with questions from past tasks stored in memory. Specifically, for each image from the current task $x^{t}$, we pair it with a question $q^{m}$ sampled from the memory. By asking past questions on new images, we prompt the model to recall and retain prior knowledge. Inspired by prior distillation-based techniques [40, 11], we employ the model from the previous task $\phi^{t-1}$ to generate answers for each new image-question pair $(x^{t},q^{m})$. These answers serve as soft pseudo-labels for the current model $\phi^{t}$, enforcing consistency with past knowledge. The pseudo-labeling loss is defined as:
$\mathcal{L}_{\text{PL}}=\mathbb{E}_{x^{t}\sim\mathcal{T}^{t}}\mathbb{E}_{q^{m}\sim\mathcal{M}}\mathcal{L}_{\text{CE}}\left[\phi^{t}(x^{t},q^{m}),\phi^{t-1}(x^{t},q^{m})\right].$ (4)
@@
We use the network’s output as soft pseudo-labels, avoiding the application of the argmax operator. Applying argmax would compel the network to align its predictions solely with the most likely class, while soft pseudo-labeling enables a more nuanced alignment with the output distribution of the model $\phi^{t-1}$ [26, 84, 66].
Our new image-question pairs $(x^{t},q^{m})$ exposes the model to a broad set of visual-question combinations, enhancing its stability across tasks. Importantly, it allows the model to retain a broad ability to answer different types of questions, even if it cannot produce precise, correct answers. Crucially, without this pseudo-labelling regularisation, the model becomes susceptible to the out-of-answer-set problem, where overfitting to the current task’s answer space leads to incorrect responses for questions from previous tasks. For instance, after training on a color recognition task, the model might incorrectly respond with a color name when asked a counting question from a prior task.
@@
Question Selection for QUAD. Randomly pairing images from the current task $x^{t}$ with past questions $q^{m}$ from memory can help mitigate forgetting, but it is inherently suboptimal as it often results in mismatched pairs. For instance, if the current macro-task $\mathcal{T}^{t}$ focuses on counting, random pairing could lead to questions like “How many cows are in this image?” being associated with an image of cars, which undermines the relevance and effectiveness of the training process. To address this, we introduce a question selection strategy that prioritises questions directly related to the object categories $c_{i}$ being learnt in the current visually-driven subtask $\mathcal{S}^{l}$. This targeted pairing ensures that the training process leverages meaningful and contextually relevant image-question combinations. By aligning questions with the visual context, QUAD enhances adaptation to new linguistically-driven tasks and visually-driven subtasks, leading to effective knowledge transfer.
@@
Example: Suppose the current visually-driven subtask $\mathcal{S}^{t}$ involves learning to count cars. Thus, the question selection strategy prioritises questions from memory relevant to cars, such as “What’s the color of the car?”, ensuring alignment between the visual content and the linguistic query.

### v2

# Ask and Remember: A Questions-Only Replay Strategy for Continual Visual Question Answering

@@ Abstract
Continual Learning in Visual Question Answering (VQACL) requires models to acquire new visual-linguistic skills (plasticity) while preserving previously learned knowledge (stability). The inherent multimodality of VQACL exacerbates this challenge, as models must balance stability across visual and textual domains while adapting to novel objects and reasoning tasks. Existing methods, primarily designed for unimodal settings, often fall short in addressing this dual requirement. In this work, we present QUestion-only replay with Attention Distillation (QUAD), a novel approach for VQACL that leverages only past task questions for regularization. By eliminating the need to store visual data, QUAD not only reduces memory overhead, but also alleviates privacy concerns. Our method introduces a Question-only Replay mechanism that selectively reuses prior task questions to counteract overfitting to the answer space of the current task, addressing the problem out of answer set. Complementing this, we propose Attention Consistency Distillation to enforce both intra-modal and inter-modal attention consistency across tasks, preserving essential visual-linguistic associations. Extensive experiments on VQAv2 and NExT-QA demonstrate that QUAD significantly outperforms state-of-the-art methods, achieving robust performance in continual VQA. Code is available at: https://github.com/IemProg/QUAD.  [^1] Work done during an internship at Computer Vision Center (CVC), Universitat Autonoma de Barcelona.
@@ 1 Introduction
Continual learning (CL) allows models to lean new skills while retaining prior knowledge, effectively mitigating catastrophic forgetting (CF) [57, 58]. This capability is essential in dynamic real-world settings, where models must continuously adapt to evolving data while preserving previously acquired knowledge [95]. Despite significant advances in CL, most research has focused on unimodal tasks (e.g. image classification) [81, 80, 38, 26, 56, 97]. However, real-world applications often require multimodal learning that integrates complex visual and textual reasoning. Visual Question Answering (VQA) stands as a representative multimodal task, requiring models to jointly interpret visual content and natural language. For instance, answering questions like “What color is the car on the left?” or “How many trucks are visible?” requires robust object recognition and a nuanced grasp of the linguistic query structure [95, 70, 88, 10, 83].
@@
The emerging field of Visual Question Answering Continual Learning (VQACL) [95] focuses on enabling models to improve VQA performance iteratively by learning from a sequence of tasks without catastrophic forgetting. Unlike static pretrained multimodal models [50, 44, 1] that rely on massive datasets and heavy computation, VQACL offers a more efficient and scalable alternative by allowing incremental adaptation without costly re-training [95, 61]. Research in VQACL is further driven by the broader scientific goal of developing multimodal agents that continuously improve their reasoning throughout deployment in a multimodal environments. A key challenge in VQACL is balancing stability—the preservation of past knowledge—and plasticity—the ability to learn new information—across both visual and linguistic modalities [95, 5]. This dual-modality requirement, combined with the need for generalisation ability, introduces additional complexity in learning. Models must retain visual and linguistic knowledge across tasks, while also generalizing to novel objects and unseen question types. For example, a model that masters counting vehicles should be able to transfer this counting capability to novel object categories, such as bicycles [27].
@@
To achieve this balance, continual VQA methods rely on memory-based replay, where previously seen examples are stored and revisited to reinforce prior knowledge and mitigate forgetting [9, 8, 77, 95]. In VQACL, this necessitates storing full image-question pairs, which demands substantial memory resources, often amounting to thousands of samples per task (e.g. 5000 samples in VQACL [95]). Storing visual data poses two core challenges. First, it incurs high computational and storage overhead, which can be infeasible in resource-constrained real-world applications. Second, more importantly, it also raises serious privacy concerns. Visual data often contains sensitive and personally identifiable information, especially in domains like healthcare, finance, and surveillance [74, 49], where stringent regulations such as GDPR [22, 91] govern data usage and storage. In contrast, textual data—such as questions—is typically generic and non-identifiable, thereby posing minimal privacy risks. Memory-free methods [41, 2, 93, 38], though eliminating storage and privacy concerns by avoiding data retention entirely, often deliver suboptimal performance in multimodal settings like VQACL. This tension between memory-rehearsal and memory-free approaches prompts a key question: Is it truly necessary to store visual data, or could retaining only past questions suffice to mitigate forgetting? To explore these questions, we propose a novel intermediate setting: VQACL with Question-only Rehearsal (VQACL-QR) (Fig. 1). In this framework, only questions from past tasks are stored—eliminating the need for visual data—and offering a practical, privacy-preserving solution for continual VQA.
Figure 2: Out-of-Answer-Set Problem in Sequential Finetuning. Confusion matrices compare Sequential Finetuning (left) and QUAD (right) across three sequentially trained tasks: Counting, Action, and Color (y-axis: answers set vocabulary, x-axis: predicted answers). Diagonal values indicate model predictions on the current task, while off-diagonal shifts reveal predictions on previous tasks. Sequential Finetuning exhibits, misclassifying past-task questions with responses from the latest task (e.g., counting questions answered with ‘Yes/No’). QUAD mitigates forgetting, preserving prior knowledge while still adapting to new tasks.
@@
To address VQACL-QR, we propose QUestion-only replay with Attention Distillation (QUAD), a novel replay framework that balances stability and plasticity using only past questions—no visual data required. Our approach introduces two key contributions, each targeting specific challenges in continual VQA. First, we introduce a Question-only Replay mechanism that leverages stored questions from previous tasks to regularize the current model through a strategic selection process that mitigates the out-of-answer-set problem, where models overfit to the current task’s answer space and consequently misanswer previous queries (see Fig. 2). Second, we propose Attention Consistency Distillation, a novel strategy that preserves attention patterns across tasks. It preserves intra-modal (text–text, image–image) and inter-modal (text–image) attention consistency, ensuring the model maintains focus on relevant regions even as it adapts to new information. Extensive experiments on standard VQACL benchmarks (VQAv2 and NExT-QA) demonstrate that QUAD achieves state-of-the-art performance, surpassing both memory-free approaches and rehearsal methods. Notably, QUAD surpasses prior methods that rely on image storage, demonstrating that storing only questions is sufficient to mitigate forgetting, validating the practicality of the Question-only Rehearsal (VQACL-QR) setting.
@@ 2 Related work
Visual Question Answering (VQA) is the task of responding to natural language queries by interpreting visual content [5, 60, 45]. Recent approaches leverage vision-language models (VLMs) built on transformer architectures [19, 71, 86, 43, 10] alongside pre-trained language models [88, 90]. For example, Cho et al. [11] introduced a generative transformer model that integrates visual and textual modalities for VQA. Many approaches enhance generalization by leveraging compositionality—a cornerstone of cognitive reasoning [35, 42]. For instance, Johnson et al. [33] explored the composition of visual attributes by creating a dataset designed for compositional reasoning, while Whitehead et al. [84] used contrastive learning to enhance compositionality, disentangling reasoning skills from visual concepts. Despite these advances, implicit decomposition strategies can hinder generalization, and crafting effective contrastive samples remains challenging.
Continual Learning (CL) seeks to develop frameworks that incrementally assimilate new data while mitigating catastrophic forgetting. This is a fundamental challenge for many deep learning methods due to catastrophic forgetting (CF) [57]. CL methods are broadly categorized into knowledge distillation-based approaches, which constrain mappings between successive models to prevent forgetting [46, 66, 31, 54, 16, 34]; optimisation-based, which adjust gradient updates to minimize interference with prior tasks [52, 7, 69, 89] and representation-based strategies that learn robust, adaptable features [21, 20, 18, 17]. Recently, prompt-based methods have emerged [82, 81, 78, 72], employing visual prompts with pre-trained transformers in CL scenarios [51]. However, since most techniques are designed for Class-Incremental Learning (CIL) in unimodal settings, they fall short when applied to multimodal data and overlook the critical issue of compositional generalization in VQA [95, 61].
Continual VQA. Recent studies have explored multimodal continual learning in VQA [13, 28, 61, 73]. However, prior works like Greco et al. [28] primarily analyze forgetting dynamics without proposing dedicated solutions, often overlooking the role of pre-trained models [59]. The VQACL benchmark [95] evaluated conventional continual learning approaches such as  [9, 8, 77, 95]. Other studies have investigated specific facets of continual VQA, including question diversity [28], compositionality [95] and domain adaptation [96]. In contrast, our work introduces a novel question-only replay mechanism with attention distillation for continual VQA that is both memory-efficient, and privacy-conscious, eliminating the need for stored prototypes as employed in the VQACL method [95] [^2] VQACL represents both the setting and the approach for continual learning in Visual Question Answering, as described in  [95]..
@@ 3.1 Setting Overview
In VQACL, the sequence of tasks is organized as a series of $L$ macro-tasks, each comprising $K$ sub-tasks, resulting in a total of $T=L\times K$ tasks. Each macro-task is designed to develop specific reasoning skills such as counting, color identification, or object recognition (i.e. linguistic task). For example, in a counting task, the model primarily engages with questions like “How many objects are there?” or “What number is shown?”.
@@
Figure 3: Overview of QUAD. QUAD is composed of three components that jointly promote stability and plasticity in VQACL setting. (1) Question-Only Memory ($\mathcal{M}$) stores questions from past tasks, without visual data. (2) Question-only replay ($\mathcal{L}_{\text{QR}}$) leverages answers generated by the previous model $\theta^{t-1}$ for new image-question pairs, encouraging the current model $\theta^{t}$ to retain past knowledge. (3) Attention Consistency Distillation ($\mathcal{L}_{\text{ACD}}$) aligns the self-attention maps between $\theta^{t-1}$ and $\theta^{t}$ to maintain focus on relevant visual-linguistic relationships. The task-specific loss ($\mathcal{L}_{\text{CE}}$) is applied solely to current task samples, promoting adaptation to new data.
@@ 3.2 Overview
We propose QUAD, a novel approach for VQACL-QR framework that avoids image storage by relying solely on previously encountered questions (see Fig. 3). Inspired by prior work in continual learning [47, 14], we adopt a regularisation framework in which the overall learning objective $\mathcal{L}_{\text{VQACL}}$ is composed of two main components:
$\mathcal{L}_{\text{VQACL}}=(1-\lambda)\mathcal{L}_{\text{Plasticity}}+\lambda\mathcal{L}_{\text{Stability}}.$ (1)
The plasticity term $\mathcal{L}_{\text{Plasticity}}$, drives the model to adapt to the current task $\mathcal{T}^{t}$, while the weighting factor $\lambda>0$ balances the trade-off between the two loss terms. Following common practice in VQA [95, 65, 4], we implement the plasticity loss using cross-entropy to compare the network’s prediction for an input image-question pair $(x^{t},q^{t})$, against the ground-truth answer $y^{t}$:
@@
The second loss component, the stability term $\mathcal{L}_{\text{Stability}}$, mitigates catastrophic forgetting. In the standard VQACL, where images from previous tasks are stored in memory $\mathcal{M}$, the stability term is computed analogously using cross-entropy, averaging the loss over triplets ${(x^{m},q^{m},y^{m})\in\mathcal{M}}$. However, under the VQACL-QR setting, storage of images $x^{m}$ from past tasks in the memory $\mathcal{M}$ is disallowed. To overcome this limitation, we design a novel stability loss $\mathcal{L}_{\text{Stability}}$ tailored for VQACL-QR. Our approach integrates two complementary losses—question-only replay $\mathcal{L}_{\text{QR}}$ and attention distillation $\mathcal{L}_{\text{ACD}}$—to effectively compensate for the absence of past task images, ensuring robust knowledge retention within the constraints of the VQACL-QR framework. Thus, the stability term is expressed as:
$\mathcal{L}_{\text{Stability}}=\mathcal{L}_{\text{QR}}+\mathcal{L}_{\text{ACD}}.$ (3)
Next, we describe the implementation of these loss components without storing past task images.
3.3 Question-only Replay
@@
To enhance knowledge retention in continual VQA, we propose a questions-only replay strategy that replays stored questions by pairing them with current task images to simulate past tasks. For each image from the current task $x^{t}$, we pair it with a question $q^{m}$ sampled from the memory. By combining stored questions with new images, our approach encourages the model to recall and reinforce previously acquired knowledge. Inspired by prior distillation-based works [47, 14], we use the model from the previous task, $\phi^{t-1}$ to generate answers for each new image-question pair $(x^{t},q^{m})$. The generated answers act as soft pseudo-labels for the current model $\phi^{t}$, enforcing consistency with prior knowledge. The loss is defined as:
$\mathcal{L}_{\text{QR}}=\mathbb{E}_{x^{t}\sim\mathcal{T}^{t}}\mathbb{E}_{q^{m}\sim\mathcal{M}}\mathcal{L}_{\text{CE}}\left[\phi^{t}(x^{t},q^{m}),\phi^{t-1}(x^{t},q^{m})\right].$ (4)
@@
Notably, we employ the network’s output as soft pseudo-labels without applying the argmax operator. Using argmax would force the network to align exclusively with the most probable class; in contrast, soft pseudo-labeling allows for a more nuanced alignment with the full output distribution of $\phi^{t-1}$ [30, 94, 76].
These image-question pairs $(x^{t},q^{m})$ expose the model to a wide range of visual-question combinations, thereby improving the model’s retention of prior knowledge. Importantly, this strategy enables the model to maintain versatility in answering diverse question types, even when predictions exhibit some uncertainty. Without this regularization, the model becomes vulnerable to the out-of-answer-set problem, wherein overfitting to the current task’s answer space results in erroneous responses for questions from prior tasks (see Fig. 2). For example, a model trained on color recognition might erroneously return a color name when confronted with a counting question from a previous task. Notably, this phenomenon closely resembles class recency bias in class-incremental learning (CIL), where models disproportionately favor newly learned classes over previously encountered ones  [68, 55].
@@
Question Selection for QUAD. Randomly pairing current task images $x^{t}$ with past questions $q^{m}$ from memory can partially mitigate forgetting but can produce semantically incoherent image-question pairs. For instance, if the current macro-task $\mathcal{T}^{t}$ focuses on counting, naive pairing might associate an image of cars with the question “How many cows are in this image?”, leading to incoherent supervision and diminished performance. To address this, we introduce a targeted question selection strategy that prioritizes questions related to the object categories $c_{i}$ in the current visually-driven subtask $\mathcal{S}^{l}$. This ensures that replayed examples remain contextually meaningful, reinforcing relevant knowledge while preventing misleading associations. By aligning stored questions with the current task’s visual concepts, QUAD facilitates effective adaptation across linguistic and visual subtasks, improving continual learning performance.
@@
Example: Suppose the current visually-driven subtask $\mathcal{S}^{t}$ involves learning to count cars. The question selection strategy prioritizes memory questions relevant to cars, such as “What’s the color of the car?”, ensuring that replay maintains associations between visual and linguistic queries.